# m-KAILIN: Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

### v1

@@ Abstract
The rapid progress of large language models (LLMs) in biomedical research has underscored the limitations of existing open-source annotated scientific corpora, which are often insufficient in quantity and quality. Addressing the challenge posed by the complex hierarchy of biomedical knowledge, we propose a knowledge-driven, multi-agent framework for scientific corpus distillation tailored for LLM training in the biomedical domain. Central to our approach is a collaborative multi-agent architecture, where specialized agents—each guided by the Medical Subject Headings (MeSH) hierarchy—work in concert to autonomously extract, synthesize, and self-evaluate high-quality textual data from vast scientific literature. These agents collectively generate and refine domain-specific question-answer pairs, ensuring comprehensive coverage and consistency with biomedical ontologies while minimizing manual involvement. Extensive experimental results show that language models trained on our multi-agent distilled datasets achieve notable improvements in biomedical question-answering tasks, outperforming both strong life sciences LLM baselines and advanced proprietary models. Notably, our AI-Ready dataset enables Llama3-70B to surpass GPT-4 with MedPrompt and Med-PaLM-2, despite their larger scale. Detailed ablation studies and case analyses further validate the effectiveness and synergy of each agent within the framework, highlighting the potential of multi-agent collaboration in biomedical LLM training. [^1] The code supporting the findings of this study is available at DropBox.
keywords:
Agentic AI, Large Language Model, Biomedical Corpus Generation
@@ 1 Introduction
The advent of large language models (LLMs) has propelled bioinformatics into a new era, enabling the development of automated solutions across a spectrum of biomedical domains [1, 2], and has demonstrated notable success for real-world biomedical question answering (QA) tasks [3]. However, the intricate and specialized nature of biomedical tasks means that general-purpose LLMs often fall short unless meticulously adapted and fine-tuned for the domain [4]. Progress in this area is further constrained by the scarcity of sufficiently large and high-quality biomedical corpora [5, 6]. While existing open-source biomedical datasets are typically of high quality, their limited scale and narrow topical coverage restrict their utility for comprehensive LLM training. Conversely, directly leveraging the expansive body of domain-specific scientific literature offers the potential to cover a broader range of biomedical topics. Nevertheless, most of these resources are unannotated, and their inherent lack of structure, coupled with the absence of QA-format organization, greatly hinders their effective use for training question-answering models. Although the vast body of biomedical literature represents a valuable and authoritative resource, its complex terminology and dense conceptual structures pose significant barriers to automated processing and dataset construction [7, 8]. Figure 1 visualizes the resulting data bottleneck. These challenges raise a crucial question: How can we automatically distill large-scale, high-quality QA pairs from scientific literature to empower biomedical LLM?
Figure 1: Analyzing the limitations and challenges of the existing pipeline.
@@
Prior research can be categorized into three primary strategies to tackle the automation of biomedical corpus distillation. First, rule-based methods[9, 10, 11] rely on human-crafted standards for data cleaning and curation, which, while effective in reducing noise, are resource-intensive and difficult to scale. Second, knowledge graph-based approaches[12, 13] structure biomedical information from texts into comprehensive graphs, but their dependence on manually curated sources limits both efficiency and scalability. Third, synthetic data generation methods[14, 15, 16, 17] use LLMs to automate the creation of QA pairs and process large corpora, yet they often lack mechanisms for interdisciplinary collaboration[18, 19], resulting in insufficient diversity and robustness in the distilled data. Recent efforts [20] have introduced a knowledge hierarchy-guided [21] approach that leverages a single LLM agent to generate and evaluate biomedical QA data with improved alignment to domain ontologies. Its reliance on a single-agent, rule-based architecture presents inherent limitations in terms of diversity, cross-domain expertise, and collaborative reasoning [22].
@@
In response to these limitations, we introduce the Multi-agent enhanced Knowledge hierArchy guIded biomedical dataset distiLlatIoN (m-KAILIN) framework—a novel, fully automated, multi-agent system designed to extract high-quality training corpora for biomedical LLMs. The m-KAILIN framework leverages a collaborative architecture of specialized LLM agents, each guided by a structured biomedical knowledge hierarchy, to evaluate and generate ‘Question-Answer-Context’ triples with strong domain alignment. The workflow begins with fine-tuned LLM agents generating candidate questions from limited annotated datasets. These questions are then matched to relevant contexts retrieved from over 23 million biomedical research articles. The question-context pair that best aligns with the knowledge hierarchy is chosen through a multi-agent evaluation and selection process. This automated pipeline facilitates the generation of high-quality, preference-based datasets that can further train LLMs to produce questions and answers aligned with biomedical ontologies, culminating in a robust, AI-ready corpus.
In summary, our principal contributions are as follows:
@@ 2 Methods
In this section, we detail the m-KAILIN framework, which leverages a collaborative multi-agent design to distill biomedical corpora from large-scale literature systematically.
@@ 2.2 Context Retrieval Agent
The Context Retrieval Agent retrieves relevant supporting documents from the large-scale PubMed [^2] PubMed: https://pubmed.ncbi.nlm.nih.gov/ corpus for each generated question. This agent employs Dense Passage Retrieval (DPR) [24], in line with Retrieval-Augmented Generation (RAG) [25] frameworks. For each question $q_{i}$, we encode both the question and every candidate document $r$ as dense vectors, $\mathbf{e}(q_{i})$ and $\mathbf{e}(r)$, using a shared BiomedBERTbase encoder [^3] BiomedBERT: https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext. The relevance score is computed as:
$p_{\eta}(r|q_{i})\propto\exp\left(\mathbf{e}(r)^{\top}\mathbf{e}(q_{i})\right),$ (3)
Documents with top-$k$ relevance scores are selected as contexts $c_{i}$ for subsequent evaluation. This decoupled retrieval enables independent optimization and clearer agent collaboration in the overall framework.
@@ 2.3 Question Evaluation Agent
The Question Evaluation Agent is designed to assess which of two candidate question-context pairs, $(q^{1}_{i},c^{1}_{i})$ and $(q^{2}_{i},c^{2}_{i})$, generated from the same source document $d_{i}$, better aligns with the hierarchical biomedical knowledge embodied in Medical Subject Headings [26] (MeSH). The agent employs a two-stage approach: an initial rule-based evaluation leveraging the MeSH ontology to create high-quality preference labels, followed by learning an automatic evaluator via LLM fine-tuning.
2.3.1 Knowledge-Guided Fine-Tuning
To enable evaluation in the absence of human-annotated preference data, we propose a knowledge-guided, rule-based cold-start method to compare candidate question-context pairs. This approach leverages the structure and information content of the MeSH ontology to provide high-confidence preference labels at scale.
Given a source document $d_{i}$ and two candidate pairs $(q^{1}_{i},c^{1}_{i})$ and $(q^{2}_{i},c^{2}_{i})$, we extract MeSH terms from $d_{i}$, $c^{1}_{i}$, and $c^{2}_{i}$. For each context $c^{j}_{i}$ ($j\in\{1,2\}$), the knowledge hierarchy similarity to $d_{i}$ is computed as:
$\bar{S}^{j}_{i}=\frac{1}{|d_{i}|\,|c^{j}_{i}|}\sum_{m^{x}\in d_{i}}\sum_{m^{y}\in c^{j}_{i}}\frac{2\times IC(\Lambda(m^{x},m^{y}))}{IC(m^{x})+IC(m^{y})},$ (4)
where $IC(m)$ is the information content for MeSH term $m$, and $\Lambda(m^{x},m^{y})$ is their lowest common ancestor. The candidate pair with the higher similarity score is labeled as the preferred sample:
$y_{i}=\arg\max_{j\in\{1,2\}}\bar{S}^{j}_{i}$ (5)
These cold-start preference labels constitute a large-scale, high-quality supervision signal for bootstrapping the subsequent LLM-based evaluation agent.
@@ 2.3.2 Evaluation Agent Preference Learning
With the knowledge-guided cold-start preference labels as supervision, we further train a large language model $\phi$ to serve as an automatic evaluation agent. Given a training set of triplets $\{(d_{i},(q^{1}_{i},c^{1}_{i}),(q^{2}_{i},c^{2}_{i}),y_{i})\}$, where $y_{i}$ indicates which candidate pair is preferred according to the MeSH-based scoring, the LLM is optimized to predict the correct preference. Formally, let $P_{\phi}(y_{i}|d_{i},q^{1}_{i},c^{1}_{i},q^{2}_{i},c^{2}_{i})$ denote the predicted probability that the $y_{i}$-th pair is preferred. The LLM is trained by minimizing the negative log-likelihood:
$\mathcal{L}_{\mathrm{eval}}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log P_{\phi}(y_{i}|d_{i},q^{1}_{i},c^{1}_{i},q^{2}_{i},c^{2}_{i}),$ (6)
@@ 2.3.3 Automatic Evaluation and Deployment
Once fine-tuned, the LLM-based evaluation agent can be deployed to automatically assess the relative quality of any two candidate question-context pairs generated from the same document. Given $(d_{i},(q^{1}_{i},c^{1}_{i}),(q^{2}_{i},c^{2}_{i}))$, the agent predicts which pair is better aligned with biomedical knowledge, thus supporting scalable and efficient data distillation for downstream tasks. This learned evaluation mechanism inherits the domain expertise embedded in the MeSH-based rules and generalizes to more complex or subtle cases, enabling robust multi-agent collaboration in the m-KAILIN framework.
@@ 2.4 Answer Generation Agent
Once the optimal question-context pair is identified, the Answer Generation Agent $\varphi$ generates high-quality answers using advanced LLMs such as GPT-4o [27]. For each input $(q_{j},c_{j})$, this agent produces an answer $\varphi(q_{j},c_{j})\rightarrow a_{j}$, facilitating the construction of comprehensive QA datasets for downstream training.
2.5 Multi-Agent Collaborative Framework
@@
This stage integrates outputs from the multiple agents to construct preference and ideal datasets, and further optimizes the question generator through direct preference optimization [28] (DPO):
@@
Step-1 Initialize Two Distinct Question Generation Agents: m-KAILIN framework will first initialize two Question Generation Agents. Specifically, one of them, $\theta^{d}$, will be based on a domain-specific LLM, e.g., BioMistral [9]. And the other $\theta^{b}$ will be based on a more general yet powerful LLM, e.g., Llama-3 [29]. Each of them will follow the same fine-tuning step introduced in Section 2.1.1 and be able to generate quality questions. We want to highlight that the reason we adopt two distinct Question Generation Agents is to enhance diversity and robustness in question generation. The domain-specific agent (e.g., BioMistral) is expected to capture fine-grained biomedical knowledge and terminology, ensuring that generated questions are highly relevant to the biomedical context. Meanwhile, the general-purpose agent (e.g., Llama-3) possesses broader reasoning and linguistic capabilities and may generate questions from more diverse perspectives. By leveraging both agents, m-KAILIN can comprehensively explore the question space over biomedical documents, increasing the likelihood of producing high-quality, informative, and non-redundant question candidates. The outputs from both agents will be used in the subsequent context retrieval and evaluation steps, allowing the downstream modules to select the most suitable question-context pairs for the construction of high-quality training datasets.
@@
Step-2 Preference Dataset Construction: By utilizing the Evaluation Agent for each generated pair from two Question Generation Agents, we select the question-context pair with the better alignment as the positive sample ($q^{+}_{i}$), and the other as the negative ($q^{-}_{i}$), forming a preference dataset $\mathcal{P}$. This preference dataset forms the supervision signal for optimizing the question generation agent in the next step.

### v2

# m-KAILIN: A Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

@@ Abstract
Corpus distillation for biomedical large language models (LLMs) seeks to address the pressing challenge of insufficient quantity and quality in open-source annotated scientific corpora, which remains a bottleneck for effective LLM training in biomedical research. This paper proposes a knowledge-driven, agentic framework for scientific corpus distillation, tailored explicitly for LLM training in the biomedical domain, addressing the challenge posed by the complex hierarchy of biomedical knowledge. Central to our approach is a collaborative multi-agent architecture, where specialized agents—each guided by the Medical Subject Headings (MeSH) hierarchy—work in concert to autonomously extract, synthesize, and self-evaluate high-quality textual data from vast scientific literature. This agentic framework collectively generates and refines domain-specific question-answer pairs, ensuring comprehensive coverage and consistency with biomedical ontologies while minimizing manual involvement. Extensive experimental results show that language models trained on our multi-agent distilled datasets achieve notable improvements in biomedical question-answering tasks, outperforming both strong life sciences LLM baselines and advanced proprietary models. Notably, our AI-Ready dataset enables Llama3-70B to surpass GPT-4 with MedPrompt and Med-PaLM-2, despite their larger scale. Detailed ablation studies and case analyses further validate the effectiveness and synergy of each agent within the framework, highlighting the potential of multi-agent collaboration in biomedical LLM training.
{IEEEkeywords}
Biomedical large language models, corpus distillation, question-answer, agentic AI
@@ 1 Introduction
The advent of large language models has propelled bioinformatics into a new era, enabling the development of automated solutions across a spectrum of biomedical domains [1, 2, 3], and has demonstrated notable success for real-world biomedical question answering (QA) tasks [4, 5]. However, the intricate and specialized nature of biomedical tasks means that general-purpose LLMs often fall short unless meticulously adapted and fine-tuned for the domain [6, 7, 8]. Progress in this area is further constrained by the scarcity of sufficiently large and high-quality biomedical corpora [9, 10, 11]. While existing open-source biomedical datasets are typically of high quality, their limited scale and narrow topical coverage restrict their utility for comprehensive LLM training. Conversely, directly leveraging the expansive body of domain-specific scientific literature offers the potential to cover a broader range of biomedical topics. Nevertheless, most of these resources are unannotated, and their inherent lack of structure, coupled with the absence of QA-format organization, greatly hinders their effective use for training question-answering models. Although the vast body of biomedical literature represents a valuable and authoritative resource, its complex terminology and dense conceptual structures pose significant barriers to automated processing and dataset construction [12, 13]. Figure 1 visualizes the resulting data bottleneck. These challenges raise a crucial question: How can we automatically distill large-scale, high-quality QA pairs from scientific literature to empower biomedical LLM?
Figure 1: Analyzing the limitations and challenges of the existing pipeline. The motivation of this study is to utilize the high-quality but limited annotated corpus to generate large-scale training corpora from raw scientific documents.
@@
Prior research can be categorized into three mainstream categories. First, rule-based methods[14, 15, 16] rely on human-crafted standards for data cleaning and curation, which, while effective in reducing noise, are resource-intensive and difficult to scale. Second, knowledge graph-based approaches[17, 18] structure biomedical information from texts into comprehensive graphs, but their dependence on manually curated sources limits both efficiency and scalability. Third, synthetic data generation methods[19, 20, 21, 22, 23] use LLMs to automate the creation of QA pairs and process large corpora. Yet, they often lack mechanisms for interdisciplinary collaboration[24, 25], resulting in insufficient diversity and robustness in the distilled data. Recent efforts [26] have introduced a knowledge hierarchy-guided [27] approach that leverages a single LLM agent to generate and evaluate biomedical QA data with improved alignment to domain ontologies. Its reliance on a single-agent, rule-based architecture presents inherent limitations in terms of diversity, cross-domain expertise, and collaborative reasoning [28].
@@
In response to these limitations, we introduce the Multi-agent enhanced Knowledge hierArchy guIded biomedical dataset distiLlatIoN (m-KAILIN) framework—a novel, fully automated, agentic framework designed to extract high-quality training corpora for biomedical LLMs. The m-KAILIN framework leverages a collaborative architecture of specialized LLM agents, each guided by a structured biomedical knowledge hierarchy, to evaluate and generate ‘Question-Answer-Context’ triples with strong domain alignment. The workflow begins with fine-tuned LLM agents generating candidate questions from limited annotated datasets. These questions are then matched to relevant contexts retrieved from over 23 million biomedical research articles. The question-context pair that best aligns with the knowledge hierarchy is chosen through a multi-agent evaluation and selection process. This automated pipeline facilitates the generation of high-quality, preference-based datasets that can further train LLMs to produce questions and answers aligned with biomedical ontologies, culminating in a robust, AI-ready corpus.
In summary, our contributions are listed as follows:
@@
2 Related Work
2.1 Dataset Distillation.
2.2 Synthesis Text Data in Biomedical field.
2.3 LLMs in Biomedical Domain.
Recent advancements [43, 44] in biomedical language models have focused on improving domain adaptation to better handle dense terminology and complex medical concepts. Approaches such as retrieval-augmented generation have enhanced models’ ability to retrieve relevant information for domain-specific tasks[45, 46, 47]. PMC-LLaMA[17] has been fine-tuned using millions of biomedical papers and medical textbooks, integrating medical knowledge graphs like UMLS for domain knowledge. Similarly, BioMistral[14] has applied extensive data processing and adaptive pretraining on PubMed Central, while HEAL[48] has combined public and proprietary datasets for continuous pretraining on a general-purpose model, further enhancing its biomedical capabilities.
@@ 3 Methods
In this section, we detail the m-KAILIN framework, which leverages an agentic framework design to distill biomedical corpora from large-scale literature systematically.
@@ 3.2 Context Retrieval Agent
The Context Retrieval Agent retrieves relevant supporting documents from the large-scale PubMed [^1] PubMed: Link corpus for each generated question. This agent employs Dense Passage Retrieval (DPR) [50], in line with Retrieval-Augmented Generation (RAG) [51] frameworks. For each question $q_{i}$, we encode both the question and every candidate document $c$ as dense vectors, $\mathbf{e}(q_{i})$ and $\mathbf{e}(c)$, using a shared BiomedBERTbase encoder [^2] BiomedBERT: Link. The relevance score is computed as:
$p(c|q_{i})\propto\exp\left(\mathbf{e}(c)^{\top}\mathbf{e}(q_{i})\right),$ (3)
Documents with top-$k$ relevance scores are selected as contexts $\textbf{c}_{i}=\{c_{1},c_{2},...,c_{k}\}$ for subsequent evaluation. This decoupled retrieval enables independent optimization and clearer agent collaboration in the overall framework.
@@ 3.3 Question Evaluation Agent
The Question Evaluation Agent is designed to assess which of two candidate question-context pairs, $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, generated from the same source document $d_{i}$, better aligns with the hierarchical biomedical knowledge embodied in Medical Subject Headings [52] (MeSH). The agent employs a two-stage approach: an initial rule-based evaluation leveraging the MeSH ontology to create high-quality preference labels, followed by learning an automatic evaluator via LLM fine-tuning.
3.3.1 Knowledge-Driven Fine-Tuning
To enable evaluation without relying on human-annotated preference data, we propose a knowledge-guided, rule-based cold-start method to compare candidate question-context pairs. This approach leverages the structural and semantic information from the MeSH ontology to generate high-confidence preference labels at scale. Given a source document $d_{i}$ and two candidate pairs $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, we can obtain MeSH terms [^3] Our dataset comprises research articles from PubMed, each accompanied by its associated MeSH terms for both the document and its context. from $d_{i}$, $\textbf{c}^{a}_{i}$, and $\textbf{c}^{b}_{i}$. Referring to Lin’s approach [53], for given context set $\textbf{c}_{i}$, the knowledge hierarchy similarity to $d_{i}$ is computed as:
$\bar{S}^{j}_{i}=\frac{1}{|d_{i}|\,|\textbf{c}_{i}|}\sum_{m^{x}\in d_{i}}\sum_{m^{y}\in\textbf{c}_{i}}\frac{2\times IC(\Lambda(m^{x},m^{y}))}{IC(m^{x})+IC(m^{y})},$ (4)
where $IC(m)$ is the information content for MeSH term $m$, and $\Lambda(m^{x},m^{y})$ is their lowest common ancestor. For any structured subject term $m$, we first calculate its information content as:
$IC(m)=-\log(\frac{freq(\mathcal{M}(m))}{n_{terms}}),$ (5)
where $\mathcal{M}(m)$ denotes the set of all descendants of MeSH term $m$, and $n_{terms}$ represents the total number of MeSH terms in the corpus. This score reflects how well the candidate context’s MeSH terms align with those of the source document. A higher $\bar{S}^{j}_{i}$ indicates that the candidate context is more semantically consistent with the document’s knowledge hierarchy. Once the similarity scores for each candidate pair have been calculated, we select the one with the higher score:
$y_{i}=\arg\max_{j\in\{a,b\}}\bar{S}^{j}_{i}$ (6)
In other words, the candidate pair whose context set exhibits a stronger semantic match to the source document is preferred. These cold-start preference labels, obtained in a rule-based and knowledge-guided manner, provide a large-scale and high-quality supervision signal. This signal can then be used to bootstrap a subsequent LLM-based evaluation agent.
@@ 3.3.2 Evaluation Agent Preference Learning
With the knowledge-guided cold-start preference labels as supervision, we further train a large language model $\phi$ to serve as an automatic evaluation agent. Given a training set of triplets $\{(d_{i},(q^{a}_{i},\textbf{c}^{a}_{i}),(q^{b}_{i},\textbf{c}^{b}_{i}),y_{i})\}$, where $y_{i}$ indicates which candidate pair is preferred according to the MeSH-based scoring, the LLM is optimized to predict the correct preference. Formally, let $P_{\phi}(y_{i}|d_{i},q^{a}_{i},\textbf{c}^{a}_{i},q^{b}_{i},\textbf{c}^{b}_{i})$ denote the predicted probability that the $y_{i}$-th pair is preferred. The LLM is trained by minimizing the negative log-likelihood:
$\mathcal{L}_{\mathrm{eval}}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log P_{\phi}(y_{i}|d_{i},q^{a}_{i},\textbf{c}^{a}_{i},q^{b}_{i},\textbf{c}^{b}_{i}),$ (7)
@@ 3.3.3 Automatic Evaluation and Deployment
Once fine-tuned, the LLM-based evaluation agent can be deployed to automatically assess the relative quality of any two candidate question-context pairs generated from the same document. Given $(d_{i},(q^{a}_{i},c^{a}_{i}),(q^{b}_{i},c^{b}_{i}))$, the agent predicts which pair is better aligned with biomedical knowledge, thus supporting scalable and efficient data distillation for downstream tasks. This learned evaluation mechanism inherits the domain expertise embedded in the MeSH-based rules and generalizes to more complex or subtle cases, enabling robust multi-agent collaboration in the m-KAILIN framework.
@@ 3.4 Answer Generation Agent
Once the optimal question-context pair is identified, the Answer Generation Agent $\varphi$ generates high-quality answers using advanced LLMs such as GPT-4o [54]. For each input $(q_{j},\textbf{c}_{j})$, this agent produces an answer $\varphi(q_{j},\textbf{c}_{j})\rightarrow a_{j}$, facilitating the construction of comprehensive QA datasets for downstream training.
3.5 Agentic Collaborative Framework
@@
This stage integrates outputs from the multiple agents to construct datasets, and further optimizes the question generator through direct preference optimization [55] (DPO):
Figure 2: The two distinct Question Generation Agents will generate different question by given the raw documents.
3.5.1 Initialize Two Distinct Question Generation Agents
@@
As depicted in Figure 2, m-KAILIN framework will first initialize two Question Generation Agents. Specifically, one of them, $\theta^{a}$, will be based on a domain-specific LLM, e.g., BioMistral [14]. And the other $\theta^{b}$ will be based on a more general yet powerful LLM, e.g., Llama-3 [56]. Each of them will follow the same fine-tuning step introduced in Section 3.1.1 and be able to generate quality questions. We want to highlight that the reason we adopt two distinct Question Generation Agents is to enhance diversity and robustness in question generation. The domain-specific agent (e.g., BioMistral) is expected to capture fine-grained biomedical knowledge and terminology, ensuring that generated questions are highly relevant to the biomedical context. Meanwhile, the general-purpose agent (e.g., Llama-3) possesses broader reasoning and linguistic capabilities and may generate questions from more diverse perspectives. By leveraging both agents, m-KAILIN can comprehensively explore the question space over biomedical documents, increasing the likelihood of producing high-quality, informative, and non-redundant question candidates.
Figure 3: The generated question pairs will be evaluated based on how well their retrieved contexts align with the knowledge hierarchy of the raw document.
3.5.2 Retrieve Context of Generated Question
The outputs from both agents will be used in the subsequent context retrieval and evaluation steps, allowing the downstream modules to select the most suitable question-context pairs for the construction of high-quality training datasets. Specifically, each pair of generated questions will feed into the Context Retrieval Agent to obtain their top-$k$ relevant documents as supplementary context (as depicted in the top region of Figure 3).
3.5.3 Preference Dataset Construction
@@
By utilizing the Evaluation Agent for each generated pair from two Question Generation Agents, we select the question-context pair with the better alignment as the positive sample ($q^{+}_{i}$), and the other as the negative ($q^{-}_{i}$), forming a preference dataset $\mathcal{P}$ (highlighted by the red dashed line in Figure 3). This preference dataset forms the supervision signal for optimizing the question generation agent in the next step.
3.5.4 Direct Preference Optimization